# AdaCtrl: Towards Adaptive and Controllable Reasoning via Difficulty-Aware Budgeting

### v1

@@ 4.1 Experimental Setup
Model and Datasets.
We adopt Qwen2.5-7B-Instruct (Qwen et al., 2025) as the backbone model for training. During the cold-start fine-tuning phase, the training problems are drawn from the DeepMATH dataset (He et al., 2025), which assigns each problem a difficulty level ranging from 1 to 9. We categorize problems with difficulty levels of 5 or below as easy, and those above 5 as hard. For easy problems, Qwen2.5-7B-Instruct is employed to generate concise responses. In contrast, for hard problems, we incorporate extensive long-form reasoning trajectories generated by Deepseek R1 (DeepSeek-AI et al., 2025). After filtering these trajectories using ground-truth answers, we construct a fine-tuning dataset comprising 8K instances, which includes 4K with short and 4K with long reasoning chains. To support difficulty-aware reinforcement learning, we further sample an additional 30K examples from DeepMATH dataset that comprise 10K easy and 20K hard problems and are distinct from those in the cold-start fine-tuning dataset.
@@
For cold-start fine-tuning, we employ the ms-swift framework (Zhao et al., 2024) for training, using a learning rate of 1e-5 and a batch size of 8. For difficulty-aware reinforcement learning, we adopt VeRL (Sheng et al., 2025), with all reinforcement learning experiments conducted under a unified setting. Specifically, we follow DAPO (Yu et al., 2025a) to eliminate KL divergence. The policy model is optimized using the AdamW optimizer with a learning rate of 1e-6, a batch size of 256, and a micro-batch size of 32. And we set the value of $\alpha$ and $\beta$ both as 0.5 during training. During the rollout phase, 16 responses are sampled per prompt, and the maximum generation length is set to 24K tokens. All experiments are conducted on NVIDIA H800 GPUs.
@@
We evaluate our method on four mathematical datasets that span both easy and challenging problems: AIME2024 (Art of Problem Solving, n.d.), AIME2025 (Art of Problem Solving, n.d.), MATH500 (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021). The first two datasets consist of more challenging Math Olympiad-style problems, whereas the latter two primarily contain simpler, grade-school level problems, with GSM8K being the easiest among them. Since AIME2024 and AIME2025 each contain only 30 samples, we report the average performance over 8 independent runs for these two datasets. All evaluations are conducted using consistent inference hyper-parameters set to a temperature of 0.7 and a top-p value of 0.8. To quantify the effectiveness and efficiency of models, we report two metrics: accuracy (Acc.) and the number of tokens generated in the response (Len.).
Baselines.
To assess the effectiveness of AdaCtrl, we compare it against several baseline methods that share the same backbone architecture. These include: (1) Qwen2.5-7B-Instruct: the unmodified base instruction-tuned model; (2) S1.1-7B: a reasoning-enhanced model obtained by further training Qwen2.5-7B-Instruct on the S1-1.1K dataset (Muennighoff et al., 2025); (3) R1-SFT: a baseline model fine-tuned Qwen2.5-7B-Instruct on a curated dataset provided by Deepseek R1, which includes problems sourced from the cold-start SFT dataset along with all responses generated by Deepseek R1; (4) R1-SFT-RL: a reinforcement learning variant of (3), trained with outcome accuracy rewards on the aforementioned 30K RL dataset; (5) Cold-Start: a model fine-tuned Qwen2.5-7B-Instruct on the constructed cold-start SFT dataset; (6) Cold-Start-RL: the reinforcement learning counterpart of (5), trained using outcome accuracy rewards on the same 30K RL dataset.
Table 1: Main results. We compute two metrics: Acc. is average accuracy and Len. is the average generated tokens of response.
@@
R1-SFT 12.08 20767.36 13.33 20184.36 62.00 8103.41 87.34 3336.81
R1-SFT-RL 21.25 18778.92 17.50 17924.35 66.80 8421.57 88.93 3896.76
Cold-Start 11.25 5553.00 7.92 6237.49 71.00 805.78 90.22 345.44
Cold-Start-RL 18.33 16911.63 14.58 15941.76 73.00 3797.60 90.67 369.94
[…]
@@ 4.2 Main Results
AdaCtrl effectively balances performance and reasoning budget. Compared to the RL baseline R1-SFT-RL, AdaCtrl achieves comparable accuracy on AIME2024, and improves performance by 1.67% on AIME2025, 7.20% on MATH500, and 2.05% on GSM8K. Simultaneously, it substantially reduces response lengths by 10.06%, 12.14%, 62.05%, and 91.04% on these datasets, respectively. When compared to S1.1-7B, AdaCtrl further demonstrates gains of 4.58%, 0.84%, and 5.20% on AIME2024, AIME2025, and MATH500, respectively, and achieves competitive results on GSM8K. Corresponding reductions in response length for this comparison are 11.21%, 13.95%, 45.13%, and 80.99%. These results suggest that AdaCtrl is capable of adaptively allocating reasoning budget based on problem difficulty, thereby achieving an effective trade-off between reasoning efficiency and effectiveness.
@@
Cold-start fine-tuning offers a strong initialization for adaptive budgeting. In contrast to R1-SFT, which is trained exclusively on long-form reasoning traces from Deepseek R1, our cold-start fine-tuning approach leverages a mixture of concise and extended reasoning trajectories. This enables the model to develop more efficient reasoning strategies, leading to substantial reductions in response length by 73.26%, 69.10%, 90.06%, and 89.65% on AIME2024, AIME2025, MATH500, and GSM8K, respectively. Moreover, when further trained with rule-based RL, models initialized from cold-start checkpoint (Cold-Start-RL) demonstrate more accurate budget control compared with that initialized from R1-SFT (i.e., R1-SFT-RL), reducing response lengths by 9.94%, 11.06%, 54.91%, and 91.04% across the same datasets. These findings highlight the importance of the cold-start fine-tuning stage in providing a solid foundation for subsequent adaptive budgeting.
@@
Designed reward functions enhance reasoning effectiveness. Using the cold-start fine-tuning model as backbone, AdaCtrl surpasses Cold-Start-RL that employs reinforcement learning based solely on outcome accuracy, achieving accuracy gains of 2.92%, 4.59%, 1.00%, and 0.31% on AIME2024, AIME2025, MATH500, and GSM8K, respectively. These improvements can be attributed to our reward design, which incorporates both difficulty estimation calibration and difficulty-aware length adjustments. These additional reward signals promote more accurate assessments of problem complexity and allocation of computational resources, leading to a more nuanced balance between effective and efficiency.
@@ 4.3 Analysis
Table 2: The comparison between adaptive and controlled reasoning modes of AdaCtrl.
AdaCtrl AIME2024 AIME2025 MATH500 GSM8K
Acc. Len. Acc. Len. Acc. Len. Acc. Len.
Adaptive 21.25 16889.50 19.17 15749.08 74.00 3195.69 90.98 349.34
[…]
Hard 21.67 17562.94 22.08 16114.87 71.20 5960.15 92.57 2058.13($\blacktriangle$ 489.15%)
4.3.1 Controllability of AdaCtrl
AdaCtrl introduces length-trigger tags that can be used as prerequisite tokens during generation to provide explicit control over response length. To evaluate the controllability of this mechanism, we simulate user intent by specifying either the “[Easy]” or “[Hard]” tag, corresponding to simplified and complex reasoning modes, respectively. The experimental results, summarized in Table 2, reveal clear trends. Compared to the adaptive reasoning mode, where the model autonomously infers the problem’s difficulty, forcing the “[Easy]” mode leads to a consistent drop in performance across all four datasets. However, it also results in a substantial reduction in response length, especially by 90.22% and 94.31% on the more challenging AIME2024 and AIME2025 datasets. In contrast, the “[Hard]” mode improves performance on most datasets and significantly increases response length, with gains of 86.51% and 489.15% on the simpler MATH500 and GSM8K datasets. These results indicate that AdaCtrl provides accurate user-level control over reasoning budgets. Furthermore, the adaptive reasoning mode demonstrates a favorable balance between effectiveness and efficiency. This suggests that our proposed framework not only enables fine-grained control but also offers enhanced flexibility to accommodate diverse requirements in practical deployments.
Figure 4: (a) The proportion of length-trigger tags across different datasets; and (b) the length of response in different levels of problems in MATH500.
4.3.2 AdaCtrl Serves as Good Difficulty Estimator
@@
To assess the problem difficulty estimation capability of AdaCtrl, we analyze the proportion of length-triggered tags generated by the model across four datasets, which serve as indicators of the model’s difficulty judgments. As shown in Figure 4 (a), AdaCtrl classifies the majority of problems in AIME2024 and AIME2025 datasets to the hard category, which include challenging math olympiad level problems. In contrast, for the MATH500 dataset, which contains a mix of easy and difficult problems (most of which are relatively solvable by current large language models), the model identifies 76.2% of the problems as easy ones. For GSM8K, the simplest dataset among the four, over 99% of the problems are categorized as easy ones. These results are consistent with the actual difficulty levels of the datasets and demonstrate that AdaCtrl develops a satisfactory capability to estimate problem difficulty through reinforcement learning.
4.3.3 AdaCtrl Facilitates Accurately Difficulty-Aware Budgeting
@@
To further investigate the adaptive difficulty-aware budgeting capabilities of AdaCtrl, we analyze its responses on the MATH500 dataset, which provides difficulty level annotations for each problem. As illustrated in Figure 4 (b), AdaCtrl generates progressively longer and more elaborate responses as the difficulty level increases from 1 to 5, ranging from approximately 0.3K to 6K tokens. This trend indicates that AdaCtrl can accurately regulate its reasoning budget based on its self-assessed estimation of problem difficulty, thereby enabling automatic and adaptive allocation of computational resources.
Figure 5: Training dynamics during RL.
4.3.4 Training Dynamics During RL
@@
To further investigate how the model learns to allocate adaptive reasoning budgets, we analyze both performance trends and budget dynamics throughout the reinforcement learning (RL) training process. As illustrated in Figure 5, the model exhibits an upward trends in performance across all four datasets, suggesting a progressive enhancement in reasoning capabilities. Regarding budget dynamics, we observe distinct patterns across datasets. On AIME2024, AIME2025, and MATH500, the average response length initially increases rapidly during early training steps, then gradually decreases and stabilizes at a level longer than the that before RL training. While for GSM8K, the response length remains relatively stable and close to that observed before RL training.

### v2

@@ 4.1 Experimental Setup
Model and Datasets. We adopt Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct (Qwen et al., 2025) as the backbone model for training. During the cold-start fine-tuning phase, the training problems are drawn from the DeepMATH dataset (He et al., 2025), which assigns each problem a difficulty level ranging from 1 to 9. We categorize problems with difficulty levels of 5 or below as easy, and those above 5 as hard. For easy problems, the backbone model is employed to generate concise responses. In contrast, for hard problems, we incorporate extensive long-form reasoning trajectories generated by Deepseek R1 (DeepSeek-AI et al., 2025). After filtering these trajectories using ground-truth answers, we construct a cold-start SFT dataset comprising 8K instances, which includes 4K with short and 4K with long reasoning chains. To support difficulty-aware reinforcement learning, we further sample an additional 30K examples from DeepMATH that comprise 10K easy and 20K hard problems and are distinct from those in the cold-start fine-tuning dataset.
Baselines.
To assess the effectiveness of AdaCtrl, we compare it against several baselines that share the same backbone model. These include: (1) Base Model: the unmodified base instruction-tuned model (i.e. Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct); (2) S1.1: a reasoning-enhanced model obtained by further training the base model on the S1-1.1K dataset (Muennighoff et al., 2025); (3) R1-SFT: a baseline model fine-tuned base model on a curated dataset that includes problems sourced from the cold-start SFT dataset along with all responses generated by Deepseek R1; (4) R1-SFT-RL: a reinforcement learning model that trained (3) with outcome accuracy rewards on the aforementioned 30K RL dataset; (5) Cold-Start: a model fine-tuned based model on the constructed cold-start SFT dataset; (6) Cold-Start-RL: a reinforcement learning model based on (5), trained using outcome accuracy rewards on the same 30K RL dataset.
@@
For cold-start fine-tuning, we employ the ms-swift framework (Zhao et al., 2024), using a learning rate of 1e-5 and a batch size of 8. For difficulty-aware reinforcement learning, we adopt VeRL (Sheng et al., 2025), with all RL experiments conducted under a unified setting. Specifically, we follow DAPO (Yu et al., 2025a) to eliminate KL divergence. The policy model is optimized using the AdamW optimizer with a learning rate of 1e-6, a batch size of 256, and a micro-batch size of 32. And we set the value of $\alpha$ and $\beta$ both as 0.5, and the value of $\delta$ as 0.625 during training. During the rollout phase, 16 responses are sampled per prompt, and the maximum generation length is set to 24K tokens. All experiments are conducted on NVIDIA H800 GPUs.
[…]
@@
We evaluate our method on four mathematical datasets that span both easy and challenging problems: AIME2024 (Art of Problem Solving, n.d.), AIME2025 (Art of Problem Solving, n.d.), MATH500 (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021). The first two datasets consist of more challenging Math Olympiad-style problems, whereas the latter two primarily contain simpler, grade-school level problems, with GSM8K being the easiest among them. Since AIME2024 and AIME2025 each contain only 30 samples, we report the average performance over 8 independent runs for these two datasets. All evaluations are conducted using consistent inference hyper-parameters set to a temperature of 0.7 and a top-p value of 0.8. In this paper, we use a baseline prompt to evaluate all baselines that do not incorporate difficulty-aware reasoning. Furthermore, we design a simple difficulty-aware prompt, derived from the baseline prompt, to support our proposed approach, as detailed in Box 4.1 for baseline prompt and Box 4.1 for difficulty-aware prompt. To quantify the effectiveness and efficiency of models, we report two metrics: accuracy (Acc.) and the number of tokens generated in the response (Len.).
Table 1: Main results of AdaCtrl’s adaptive mode. We compute two metrics: Acc.(%) is average accuracy and Len. is the average generated tokens.
@@
R1-SFT-7B 12.08 20767.36 13.33 20184.36 62.00 8103.41 87.34 3336.81
Cold-Start-SFT-7B 11.25 5553.00 7.92 6237.49 71.00 805.78 90.22 345.44
R1-SFT-RL-7B 21.25 18778.92 17.50 17924.35 66.80 8421.57 88.93 3896.76
Cold-Start-RL-7B 18.33 16911.63 14.58 15941.76 73.00 3797.60 90.67 369.94
AdaCtrl-7B 21.25 16889.50 19.17 15749.08 74.00 3195.69 90.98 349.34
Qwen2.5-14B-Instruct
Qwen2.5-14B-Instruct 11.67 1043.20 10.42 1136.26 73.60 568.28 93.86 215.42
S1.1-14B 32.92 17278.03 29.58 16167.50 76.00 4393.17 95.15 1460.26
R1-SFT-14B 19.58 20485.99 20.42 18447.68 65.80 6473.98 90.67 2883.76
Cold-Start-SFT-14B 12.92 4734.60 12.50 6431.17 76.20 929.48 93.63 287.72
R1-SFT-RL-14B 24.17 18549.26 23.33 18294.04 70.60 6397.75 91.05 2814.69
Cold-Start-RL-14B 27.50 17722.25 24.17 17088.44 75.60 4657.38 94.01 378.39
AdaCtrl-14B 34.58 15173.21 25.83 14476.93 79.20 3209.84 94.09 316.12
@@ 4.2 Main Results
AdaCtrl Effectively Balances Performance and Reasoning Budget.
@@
AdaCtrl demonstrates competitive overall performance, outperforming most baselines across four datasets while efficiently managing the reasoning budget. Specifically, compared to the RL baseline R1-SFT-RL based on Qwen2.5-7B-Instruct, AdaCtrl-7B achieves comparable accuracy on AIME2024 and further improves accuracy by 1.67% on AIME2025, 7.20% on MATH500, and 2.05% on GSM8K. At the same time, it substantially reduces response lengths by 10.06%, 12.14%, 62.05%, and 91.04% on these datasets, respectively. Under the 14B setting, AdaCtrl-14B increases accuracy by 10.41%, 2.5%, 8.6%, and 3.04% on AIME2024, AIME2025, MATH500, and GSM8K, while compressing the reasoning budget by 18.20%, 20.92%, 49.83%, and 88.77% on the corresponding datasets. Similar trends are also observed when comparing AdaCtrl to the S1.1-7B and S1.1-14B baselines. Overall, these results suggest that AdaCtrl can adaptively allocate the reasoning budget according to problem difficulty, thereby effectively balancing reasoning efficiency and effectiveness.
Cold-start Fine-tuning Provides an Effective Foundation for Adaptive Budgeting.
@@
Unlike R1-SFT, which is trained solely on long-form reasoning traces from Deepseek R1, our cold-start fine-tuning strategy (i.e., Cold-Start-SFT) incorporates a combination of both concise and extended reasoning trajectories. This diverse training regimen enables the model to acquire more efficient reasoning strategies, resulting in substantial reductions in response length by 73.26%, 69.10%, 90.06%, and 89.65% for AIME2024, AIME2025, MATH500, and GSM8K, respectively, on the 7B model, and by 76.88%, 54.14%, 85.64%, and 90.02% on the 14B model. Furthermore, when models initialized from the cold-start checkpoint undergo additional reinforcement learning (i.e., Cold-Start-RL), they demonstrate superior budget control compared to those initialized from R1-SFT (i.e.,R1-SFT-RL), yielding additional reductions in response length of 9.94%, 11.06%, 54.91%, and 91.04% across the same datasets on the 7B model, and 4.46%, 6.61%, 27.20%, and 86.56% on the 14B model. These results underscore the critical role of the cold-start fine-tuning phase in establishing a robust foundation for effective adaptive budgeting in downstream tasks.
Our Reward Design Enhance Both Reasoning Effectiveness and Efficiency.
@@
Building upon the cold-start fine-tuning model, AdaCtrl outperforms Cold-Start-RL, which applies reinforcement learning based solely on outcome accuracy by achieving notable accuracy improvements of 2.92%, 4.59%, 1.00%, and 0.31% on AIME2024, AIME2025, MATH500, and GSM8K, based on with the 7B model, and improvements of 7.08%, 1.66%, 3.6%, and 0.08% with the 14B model with much less token consumption. These accuracy gains are attributable to our reward function design, which incorporates both difficulty estimation calibration and difficulty-aware length adjustments. By introducing these additional reward signals, our approach enables the model to more accurately assess problem complexity and allocate computational resources accordingly, thereby achieving a more refined balance between reasoning effectiveness and efficiency.
Table 2: Comparison of adaptive and controlled reasoning modes in AdaCtrl. The easy mode reduces reasoning tokens but at the cost of lower performance, making it suitable for fast-response scenarios. Hard mode uses more tokens and yields better results. The adaptive mode automatically estimates problem difficulty to balance effectiveness and efficiency.
Model AIME2024 AIME2025 MATH500 GSM8K
[…]
AdaCtrl-7B (Adaptive) 21.25 16889.50 19.17 15749.08 74.00 3195.69 90.98 349.34
AdaCtrl-7B (Easy) 14.58 1652.42 10.00 896.14 70.80 652.76 90.75 314.49
AdaCtrl-7B (Hard) 21.67 17562.94 22.08 16114.87 71.20 5960.15 92.57 2058.13
AdaCtrl-14B (Adaptive) 34.58 15173.21 25.83 14476.93 79.20 3209.84 94.09 316.12
AdaCtrl-14B (Easy) 15.83 1394.78 11.25 985.35 75.40 578.37 94.31 284.83
AdaCtrl-14B (Hard) 36.25 15004.55 26.25 14564.21 74.40 5321.25 93.33 2598.28
[…]
AdaCtrl-7B AIME2024 AIME2025 MATH500 GSM8K
[…]
$\alpha=0.5,\beta=0.5,\delta=0.625$ 21.25 16889.50 19.17 15749.08 74.00 3195.69 90.98 349.34
$\alpha=1.0,\beta=1.0,\delta=0.625$ 23.75 14762.78 20.00 13986.64 73.80 3134.37 90.30 325.04
$\alpha=0.5,\beta=1.0,\delta=0.625$ 22.08 15616.19 18.75 16116.17 73.80 2836.52 91.13 324.10
$\alpha=1.0,\beta=0.5,\delta=0.625$ 21.67 17843.66 22.08 16684.28 73.40 3727.77 91.66 365.03
$\alpha=0.5,\beta=0.5,\delta=0.5$ 22.92 13790.53 20.42 14040.20 74.00 2888.39 90.60 388.10
$\alpha=0.5,\beta=0.5,\delta=0.375$ 20.00 13385.01 20.00 13689.74 74.60 2181.96 90.60 323.27
Human-in-the-Loop Controllability of AdaCtrl.
AdaCtrl introduces difficulty-aware tags that serve as prerequisite tokens during generation, offering explicit control over response length. To assess the controllability of this mechanism, we emulate user intent by specifying either the “[Easy]” or “[Hard]” tag, which correspond to simplified and complex reasoning modes, respectively. As presented in Table 2, the experimental results demonstrate that our approach affords effective control over the reasoning budget. Compared to the adaptive reasoning mode, where the model autonomously determines the problem’s difficulty, enforcing the “[Easy]” mode consistently leads to reduced performance across all four datasets. However, it also achieves a substantial decrease in response length by 90.22% and 94.31% on the more challenging AIME2024 and AIME2025 datasets for the 7B model, and by 90.81% and 93.19% for the 14B model, respectively. Conversely, the “[Hard]” mode enhances performance on most datasets and markedly increases response length, with gains of 86.51% and 489.15% on the simpler MATH500 and GSM8K datasets under the 7B setting, and increases of 65.78% and 721.93% under the 14B setting. These findings indicate that AdaCtrl enables precise, human-in-the-loop control over reasoning budgets.
@@ 4.3 Analysis
Table 4: Ablation study based on 7B model. $r_{f}$ is the difficulty-estimation calibration reward and $r_{l}$ is the difficulty-aware length reward.
Acc.(%) AIME2024 AIME2025 MATH500 GSM8K
AdaCtrl-7B 21.25 19.17 74.00 90.98
w/o $r_{f}$ 17.08 (-4.17%) 16.25 (-2.92%) 72.20 (-1.80%) 89.92 (1.06%)
w/o $r_{l}$ 15.42 (-5.83%) 16.67 (-2.50%) 68.60 (-5.4%) 90.60 (-0.38%)
Hyper-parameter Analysis.
To assess the robustness of our proposed method, we conduct two groups of hyperparameter analyses, as summarized in Table 3. Specifically, we examine: (1) the weights $\alpha$ and $\beta$, which correspond to the difficulty-estimation calibration reward and the difficulty-aware length reward, respectively; and (2) the difficulty threshold $\delta$. For reward weights analysis, we investigate the effect of varying the reward weights across different combinations $\{1:0.5:0.5,1:1:1,1:0.5:1,1:1:0.5\}$. We observe that accuracy remains largely stable while output length exhibits slight variations. For example, on AIME2024, accuracy stays within 21–23% as the response length ranges from 14K to 17K tokens. Similarly, on GSM8K, accuracy consistently falls between 90–91%, while the response length varies from 325–365 tokens. For difficulty threshold analysis, we further sweep the difficulty threshold $\delta$ over $\{0.625,,0.5,,0.375\}$. A smaller value of $\delta$ categorizes more samples as Easy, leading to shorter outputs while maintaining comparable accuracy. For instance, on MATH500, the response length decreases from 3195.69 to 2888.39 and then to 2181.96 tokens, while accuracy remains around 74%. These results collectively demonstrate the robustness of our proposed approach under different hyperparameter settings.
Figure 3: The proportion dynamics of difficulty-aware tags across different datasets during reinforcement learning.
Ablation Study.
To further assess the effectiveness of the proposed reward functions, we conducted ablation studies by individually removing the difficulty estimation calibration reward $r_{f}$ and the difficulty-aware length reward $r_{l}$. As shown in Table 4, the removal of either reward leads to a noticeable decline in performance across all four datasets. These results demonstrate that the combined use of all designed rewards is essential for achieving better optimization.
Figure 4: (a) The length of response in different difficulty levels of problems in MATH500, where higher levels indicate more challenging problems; (b) Controllability comparison of AdaCtrl and prompting-based approach.
AdaCtrl Serves as Good Difficulty Estimator.
@@
To better assess the difficulty estimation capability of AdaCtrl, we analysis the proportion of difficulty-aware tags generated by AdaCtrl-7B across four datasets during reinforcement learning (RL). As illustrated in Figure 3, we first observe that, at the initial stage, the model tends to classify most samples in all datasets as easy. This is likely because the cold-start SFT primarily teach the model to explicitly generate self-aware difficulty tags in the expected format, rather than to accurately assess problem difficulty from its own perspective. However, following our designed difficulty-aware RL, AdaCtrl predominantly classifies the majority of problems in the AIME2024 and AIME2025 datasets as hard. These datasets consist of challenging math Olympiad-level problems. In contrast, in the MATH500 dataset, which contains a mixture of easy and difficult problems (the majority of which are relatively solvable by current large language models), the model identifies 76.2% of problems as easy. For GSM8K, the simplest dataset among the four, over 99% of problems are categorized as easy, which also accounts for the superior performance of AdaCtrl-14B (Easy) on GSM8K as reported in Table 2. These results align with the actual difficulty levels of the datasets and demonstrate that AdaCtrl develops a robust capability to estimate problem difficulty through RL.
AdaCtrl Facilitates Accurate Difficulty-Aware Budgeting.
@@
To further investigate the adaptive difficulty-aware budgeting capabilities of AdaCtrl, we analyze AdaCtrl-7B’s responses on the MATH500 dataset, which provides difficulty level annotations for each problem. As illustrated in Figure 4 (a), AdaCtrl generates progressively longer and more elaborate responses as the difficulty level increases from 1 to 5, ranging from approximately 0.3K to 6K tokens. This trend indicates that AdaCtrl can accurately regulate its reasoning budget based on its self-assessed estimation of problem difficulty, thereby enabling automatic and adaptive allocation of computational resources.
Controllability Analysis.
Previous approaches often rely on vanilla prompting methods to achieve token-level control over reasoning budgets. To assess the controllability of AdaCtrl, we compare it with a prompting-based baseline that uses the S1.1-7B model as the backbone and follows Nayab et al. (2025) to augment user prompts with the instruction: "Limit the length of the answer to 500 tokens." The response lengths generated by AdaCtrl in both easy and adaptive modes, as well as those produced by S1.1-7B and its prompting-based budget control variant, are reported in Figure 4 (b). We can observe that, through the prompting-based method (i.e., S1.1-7B-PE) explicitly restricting outputs to 500 tokens, it only yields reductions in response length of only 5.25%, 8.96%, 7.97%, and 12.48% on the AIME2024, AIME2025, MATH500, and GSM8K datasets, respectively, falling significantly short of the targeted 500-token limit. We also observed similar trends when the length constraint was set to 1000 tokens. These results suggest that achieving precise, fine-grained control over output length is challenging due to highly dependent on model’s instruction-following capabilities, and that current models may struggle to reliably interpret and execute such granular prompts. In contrast, the easy mode of AdaCtrl achieves substantially greater compression of reasoning budgets, reducing response lengths by 90.22%, 94.32%, 79.57%, and 9.98% across the same datasets. This demonstrates superior controllability, which can be primarily attributed to AdaCtrl’s mixed fine-tuning strategy and its design of a difficulty-aware length reward.
Figure 5: The length distribution of response generated by AdaCtrl-7B.
Length Distribution of Generated Responses.
To further demonstrate the budget allocation capability of AdaCtrl, we analyze the response length distribution of AdaCtrl-7B across all inference results from the evaluated benchmarks. Specifically, we compute the distributions separately for samples categorized as easy and hard by the model predicted tags. As shown in Figure 5, AdaCtrl clearly differentiates between easy and hard problems. Notably, the response lengths for easy problems are concentrated within a relatively narrow range. These findings indicate that our approach achieves effective and accurate budget control guided by the self-assessed difficulty-aware tags.
Figure 6: Training dynamics of accuracy and response length across different datasets during reinforcement learning.
Training Dynamics During RL.
@@
To further investigate how the model learns to allocate adaptive reasoning budgets, we analyze both performance trends and budget dynamics of AdaCtrl-7B throughout the reinforcement learning (RL) training process. As illustrated in Figure 6, the model exhibits upward trends in performance across all four datasets, suggesting a progressive enhancement in reasoning capabilities. Regarding budget dynamics, we observe distinct patterns across datasets. On AIME2024, AIME2025, and MATH500, the average response length initially increases rapidly during early training steps, then gradually decreases and stabilizes at a level longer than that before RL training. While for GSM8K, the response length remains relatively stable and close to that observed before RL training.